# Breaking Scale Anchoring: Frequency Representation Learning for Accurate High-Resolution Inference from

### v1

@@ 1 Introduction
Traditional numerical simulation methods in Spatiotemporal Forecasting (STF) can achieve high accuracy yet incur substantial computational costs (Choi & Moin, 2011). Recent research utilizing deep learning methods for STF can effectively balance accuracy and computational efficiency, showing potential to replace numerical methods (Zhang et al., 2023; Saad et al., 2024). However, high-resolution, high-fidelity Direct Numerical Simulations (DNS) that provide training data for deep learning methods face extremely high costs. Even with high-resolution data available, the extremely high resolution poses impossible training VRAM requirements for current hardware, while the requirements for inference are much smaller. Zero-shot super-resolution (ZS-SR) deep learning models can leverage low-cost, low-resolution data for low-VRAM training to perform high-resolution STF (Li et al., 2020).
@@
Existing ZS-SR STD methods are based on Neural Operators (NOs). This is because, unlike general neural network architectures that learn function mappings from point space to point space, NOs learn functional mappings from function to function (Lu et al., 2021). Since functions can be discretized into spaces of arbitrary resolution, NOs naturally meets the requirements for inputs and outputs of different resolutions (Kovachki et al., 2023). On the other hand, many variants, such as Fourier NO, learn parameters in a scale-agnostic manner in the frequency domain, which can also improve the generalization of the model’s physical laws (Li et al., 2020).
@@
Existing researchs expect models’ successful generalization to maintain similar accuracy across inputs of different resolutions (Talebi & Milanfar, 2021; Gao et al., 2025). From this perspective, our pilot study on ZS-SR STF for all mainstream architectures in Section 3 shows that every baseline exhibits excellent generalization. However, for a $p$-th order numerical solver simulating at a resolution $\alpha$ times higher than the low resolution, the error theoretically decrease by a factor of $\alpha^{p}$. This gap arises from the fact that the upper bound of the frequency of physical laws that low-resolution data can represent is limited by the Nyquist frequency of the training data. When models are trained on low-resolution data and perform inference at high resolution: the model struggles to handle unseen high-frequency components, causing the model’s error to be anchored at the low-resolution data. We refer to this data-driven limitation as Scale Anchoring, which is fundamentally different from previously studied issues such as Spectral Bias (SB) and Discretization Mismatch Error (DME).
@@
Addressing Scale Anchoring requires improving generalization to higher relative frequencies that are not present at the training grid. Prior methods for cross-resolution generalization of physical laws were not explicitly designed to tackle this scale-anchoring mechanism (Li et al., 2020; 2024c). We therefore propose Frequency Representation Learning (FRL): (i) construct multi-resolution training data via downsampling; (ii) introduce Nyquist-normalized frequency representations that yield resolution-invariant embeddings for the same physical frequencies; and (iii) add a frequency-aware loss to promote spectral consistency across scales. Steps (i) and (iii) follow standard practices in multi-scale training and spectral regularization, while Step (ii), which is aligned with Scale Anchoring, is novel. Under mild assumptions, our analysis shows that this alignment encourages more stable high-frequency bands. It allows FRL to reduce the high-resolution error of a given baseline as resolution increases, although it does not guarantee strict order convergence like numerical solvers.
@@ 2.1 Zero-Shot Super-Resolution Spatiotemporal Forecasting
STF tasks take one or multiple physical spatial field snapshots at different time steps as input, predict the next time step’s snapshot, and repeat iteratively. ZS-SR STF tasks train models on low-resolution snapshots only, but at inference time take high-resolution physical spatial field snapshots at different time steps as input, predict the next time step’s snapshot, and repeat iteratively. NOs are naturally suited for this task due to their resolution-agnostic input and output capabilities (Li et al., 2020; Kovachki et al., 2023). Existing NO variants span multiple directions: FNO pioneered ZS-SR through frequency-domain learning (Li et al., 2020; Jiang et al., 2023; Atif et al., 2024); PINO enhanced accuracy via physics constraints (Li et al., 2024c); TNO specialized temporal modeling for long-term predictions (Diab & Al-Kobaisi, 2025); and multi-scale methods like Multi-Grid Tensorized FNO and U-FNO reduced computational complexity through hierarchical decomposition (Kossaifi et al., 2023; Wen et al., 2022). While these methods improve resolution generalization, they do not explicitly address the fundamental frequency limitation imposed by the training data’s Nyquist frequency. Without mechanisms to learn or extrapolate frequency patterns beyond this hard boundary, they remain fundamentally unable to resolve Scale Anchoring.
@@
SB describes the tendency of neural networks to fit target functions from low to high frequencies within the training Nyquist band, leading to larger errors on high-frequency components that are actually present in the supervision data (Xu et al., 2019; Rahaman et al., 2019). To mitigate SB, prior work introduces Fourier feature mappings and positional encodings, periodic activation functions, multi-scale or hierarchical architectures, and explicit spectral or frequency-weighted losses, all of which strengthen the network’s ability to represent and fit high-frequency content already contained in the data (Tancik et al., 2020; Sitzmann et al., 2020; Liu et al., 2024; 2025).
In the NO literature, lack of discretization-invariance (lack of DI) and DME formalize the problem that the same operator network can produce inconsistent outputs on different grids (Bartolucci et al., 2023; Gao et al., 2025; Lanthaler et al., 2024). These works analyze aliasing and discretization error to propose alias-free spectral parameterizations, multi-grid training, and carefully designed interpolation operators to enforce cross-grid consistency. More broadly, multi-resolution generation in vision and scientific ML—whether cascaded, patch-based, or training-free—expose models to multiple resolutions (Tian et al., 2023; He et al., 2023; Ho et al., 2022; Bar-Tal et al., 2023; Yu et al., 2023). In this sense, the first and third steps of FRL (multi-resolution training and frequency-aware loss) follow standard practices in these research directions and are not methodologically novel.
@@ 4.2 Experiment Validation
The rapid decay of high-frequency components energy in Figure 3 reveals the deeper mechanism behind the phenomenon in Figure 2: as shown by Gao et al. (2025), while high-resolution single-step discretization primarily introduces errors in the low-frequency portion, a more dominant mechanism during multi-step inference is that models cannot maintain frequencies above the training Nyquist frequency. This results in the pattern shown in Table 2 where, during high-resolution inference, the proportion of low-frequency error decreases while high-frequency error becomes dominant due to more severe accumulation and amplification. These experimental results validate Theorem 2, confirming that frequency components above the training resolution’s Nyquist frequency dominate the error during high-resolution inference.
@@ 6.1.1 Datasets
Fluid Simulation: Non-reacting HIT consists of 3D homogeneous isotropic turbulence simulations with a spherical $\text{O}_{2}$ core (radius $0.25L$) embedded in $\text{CH}_{4}$ (Chung et al., 2022). The dataset tracks velocity components, thermodynamic variables, and species mass fractions across 98 timesteps spanning 34 microseconds. The temporal sequence is split into 70 timesteps for training, 14 for validation, and 14 for testing. Training resolution: $32^{3}$ grid; Test resolutions: $43^{3}$, $64^{3}$, and $129^{3}$. Low-resolution fields are obtained by uniformly coarsening the original $129^{3}$ DNS snapshots by factors of $2\times$, $3\times$, and $4\times$ in each spatial direction (i.e., $129^{3}\!\to\!64^{3},43^{3},32^{3}$) using spectral low-pass filtering followed by strided sampling. We choose $32^{3}$ as the training resolution because it is the coarsest grid that still preserves the main flow structures and allows stable training for all baselines, while maximizing the Nyquist gap between the training grid and the finest test grid.
Weather Forecasting: ERA5 contains global atmospheric reanalysis data across six vertical pressure levels (200–1000 hPa) (Hersbach et al., 2020; Copernicus Climate Change Service (2017), C3S). Physical variables include temperature $T$ (K), horizontal wind components $(u,v)$ (m/s), and geopotential height $z$ ($\text{m}^{2}/\text{s}^{2}$). The dataset spans 360 days (January 2024–January 2025) with 6-hour temporal resolution, totaling 4320 timesteps, with the final 7 days reserved for validation. Training resolution: $180\times 90\times 6$; Test resolutions: $360\times 180\times 6$, $720\times 361\times 6$, and $1440\times 721\times 6$. Here the original ERA5 fields are stored on a $1440\times 721\times 6$ grid; we construct the low-resolution training data by regridding to $180\times 90\times 6$ via area-weighted averaging.
@@ 6.1.2 Baseline Selection and Implementation Details
We selected the latest SOTA baselines for each architecture published in NeurIPS/ICLR/ICML. Fluid Simulation: Neural SPH (GNN), DeepLag (Transformer), PARCv2 (CNN), DYffusion (Diffusion), SFNO (NO), FNODE (Neural ODE), NeuralFluid (NN) (Toshev et al., 2024; Ma et al., 2024; Nguyen et al., 2024; Rühling Cachay et al., 2023; Cao et al., 2024; Li et al., 2024a; b). Weather Forecasting: WeatherGFT (Transformer), PDE-CNN (CNN), ARCI (Diffusion), Graph-EFM (GNN), ClimODE (Neural ODE) (Xu et al., 2024; Donà et al., 2020; Rühling Cachay et al., 2023; Oskarsson et al., 2024; Verma et al., 2024). FRL-enhanced baselines follow a unified training recipe; full hyperparameter settings are provided in Appendix C.
@@ 7 Discussion
Reproducibility Statement
We support reproducibility via: (i) complete algorithm pseudocode and implementation details (Appendix C); (ii) a precise experimental protocol covering datasets and splits, preprocessing, baseline selection, matched training budgets, evaluation metrics, and validation criteria (Section 6.1). For completeness and transparency, we also include computational/training complexity analysis (Appendix E) and ablation plus parameter sensitivity of selected hyperparameters (Appendix H). After the review process, we will release source code, configuration files to reproduce all tables/figures, pretrained checkpoints, data-preparation scripts, and archived commit hashes.
@@ Appendix G Complete Experimental Results
In addition to the 3D scenarios presented in Section 6.1, we conducted comprehensive experiments on 2D fluid simulation and weather forecasting tasks: (1) MegaFlow2D provides large-scale 2D external flow simulations around circular and elliptical obstacles (Xu et al., 2023). The dataset contains 2000 distinct flow configurations with approximately 900 temporal snapshots each at 0.01s intervals. We use 1600 configurations for training, 200 for validation, and 200 for testing. Physical variables include velocity fields $(u,v)$ and pressure field $p$. All simulations maintain Reynolds number $Re=300$, fluid density $\rho=1\times 10^{3}$ $\text{kg/m}^{3}$, and kinematic viscosity $\nu=1\times 10^{-3}$ $\text{m}^{2}/\text{s}$ within a $20\text{m}\times 10\text{m}$ domain. Training resolution: $64\times 32$ grid; Test resolutions: $128\times 64$, $256\times 128$, and $512\times 256$. (2) ERA5-500hPa contains global atmospheric reanalysis data at the 500 hPa pressure level (Hersbach et al., 2020; Copernicus Climate Change Service (2017), C3S). Physical variables include temperature $T$ (K), horizontal wind components $(u,v)$ (m/s), and geopotential height $z$ ($\text{m}^{2}/\text{s}^{2}$). The dataset spans 360 days with 6-hour temporal resolution. Training resolution: $180\times 90$; Test resolutions: $360\times 180$, $720\times 361$, and $1440\times 721$.
@@
To enable direct inference at arbitrary resolutions without interpolation-based resampling, we adapt all baseline architectures following three unified design principles. First, we eliminate resolution-dependent components by replacing fixed-size operations with their resolution-agnostic counterparts: fully convolutional layers that can process inputs of arbitrary spatial dimensions (Long et al., 2015), dynamic graph construction that adapts to varying node counts (Zheng et al., 2024), and variable-length token sequences in transformers that naturally handle different sequence lengths (Vaswani et al., 2017; Zhai et al., 2023). Second, we adopt normalized coordinate systems where spatial positions are mapped to $0,1]^{d}$ regardless of the actual resolution, ensuring that learned spatial relationships remain valid across scales. This coordinate-based representation approach has been successfully demonstrated in implicit neural representations (Sitzmann et al., 2020; Mildenhall et al., 2020), where networks learn continuous functions of normalized coordinates rather than discrete grid positions. Third, we condition the models on resolution information either implicitly through the coordinate normalization or explicitly through resolution embedding vectors, allowing the network to adapt its processing based on the sampling density. This scale-aware conditioning strategy follows principles established in multi-scale network architectures (Tan & Le, 2019), enabling models to adjust their computational patterns based on the input resolution. These modifications preserve each architecture’s inductive biases (convolutions maintain translation equivariance, graph networks preserve permutation invariance (Wu et al., 2020), and attention mechanisms maintain their global receptive field) while enabling deployment at resolutions beyond those seen during training. In summary, we adopted various architecture-specific strategies to achieve multi-resolution inference. Although these methods differ from each other and may not be elegant, they allow us to verify that Scale Anchoring is a universally present problem. More importantly, these methods enable us to validate that the scale anchoring phenomenon persists even when models operate directly at target resolutions. We confirm that this phenomenon stems from fundamental limitations in the training data’s frequency content, rather than architectural constraints.
@@
Furthermore, we compare against SOTA ZS-SR methods specifically designed for spatiotemporal forecasting: FNO and PINO for fluid simulation (Li et al., 2020; 2024c), and TNO and Climate FNO for weather prediction (Diab & Al-Kobaisi, 2025; Jiang et al., 2023). These methods represent the current best practices in zero-shot super-resolution for their respective domains.

### v2

@@ 1 Introduction
Traditional numerical simulation methods in Spatiotemporal Forecasting (STF) can achieve high accuracy yet incur substantial computational costs (Choi2011GridpointRF). Recent research utilizing deep learning methods for STF can effectively balance accuracy and computational efficiency (Zhang2023NowcastNet; Saad2024BayesNF). However, high-resolution, high-fidelity Direct Numerical Simulations (DNS) that provide training data for deep learning methods face extremely high costs. Even with high-resolution data available, the extremely high resolution poses impossible training VRAM requirements for current hardware, while the requirements for inference are much smaller. Zero-shot super-resolution (ZS-SR) deep learning models can leverage low-cost, low-resolution data for low-VRAM training to perform high-resolution STF (li2020fourier).
@@
Existing ZS-SR STD methods are based on Neural Operators (NOs). This is because, unlike general neural network architectures that learn function mappings from point space to point space, NOs learn functional mappings from function to function (Lu2021). Since functions can be discretized into spaces of arbitrary resolution, NOs naturally meets the requirements for inputs and outputs of different resolutions (kovachki2023neural). On the other hand, many variants, such as Fourier NO, learn parameters in a scale-agnostic manner in the frequency domain, which can also improve the generalization of the model’s physical laws (li2020fourier).
@@
Existing researchs expect models’ successful generalization to maintain similar accuracy across inputs of different resolutions (talebi2021learning; gao2025discretization). From this perspective, our pilot study on ZS-SR STF for all mainstream architectures in Section 3 shows that every baseline exhibits excellent generalization. However, for a $p$-th order numerical solver simulating at a resolution $\alpha$ times higher than the low resolution, the error theoretically decrease by a factor of $\alpha^{p}$. This gap arises from the fact that the upper bound of the frequency of physical laws that low-resolution data can represent is limited by the Nyquist frequency of the training data. When models are trained on low-resolution data and perform inference at high resolution: the model struggles to handle unseen high-frequency components, causing the model’s error to be anchored at the low-resolution data. We refer to this data-driven limitation as Scale Anchoring, which is fundamentally different from previously studied issues such as Spectral Bias (SB) and Discretization Mismatch Error (DME).
@@
Addressing Scale Anchoring requires improving generalization to higher relative frequencies that are not present at the training grid. Prior methods for cross-resolution generalization of physical laws were not explicitly designed to tackle this scale-anchoring mechanism (li2020fourier; li2024physics). We therefore propose Frequency Representation Learning (FRL): (i) construct multi-resolution training data via downsampling; (ii) introduce Nyquist-normalized frequency representations that yield resolution-invariant embeddings for the same physical frequencies; and (iii) add a frequency-aware loss to promote spectral consistency across scales. Steps (i) and (iii) follow standard practices in multi-scale training and spectral regularization, while Step (ii), which is aligned with Scale Anchoring, is novel. Under mild assumptions, our analysis shows that this alignment encourages more stable high-frequency bands. It allows FRL to reduce the high-resolution error of a given baseline as resolution increases, although it does not guarantee strict order convergence like numerical solvers.
@@ 2.1 Zero-Shot Super-Resolution Spatiotemporal Forecasting
STF tasks take one or multiple physical spatial field snapshots at different time steps as input, predict the next time step’s snapshot, and repeat iteratively. ZS-SR STF tasks train models on low-resolution snapshots only, but at inference time take high-resolution physical spatial field snapshots at different time steps as input, predict the next time step’s snapshot, and repeat iteratively. NOs are naturally suited for this task due to their resolution-agnostic input and output capabilities (li2020fourier; kovachki2023neural). Existing NO variants span multiple directions: FNO pioneered ZS-SR through frequency-domain learning (li2020fourier; jiang2023efficient; atif2024fourier); PINO enhanced accuracy via physics constraints (li2024physics); TNO specialized temporal modeling for long-term predictions (diab2025temporal); and multi-scale methods like Multi-Grid Tensorized FNO and U-FNO reduced computational complexity through hierarchical decomposition (kossaifi2023multi; wen2022u). While these methods improve resolution generalization, they do not explicitly address the fundamental frequency limitation imposed by the training data’s Nyquist frequency. Without mechanisms to learn or extrapolate frequency patterns beyond this hard boundary, they remain fundamentally unable to resolve Scale Anchoring.
@@
SB describes the tendency of neural networks to fit target functions from low to high frequencies within the training Nyquist band, leading to larger errors on high-frequency components that are actually present in the supervision data (xu2019frequency; rahaman2019spectral). To mitigate SB, prior work introduces Fourier feature mappings and positional encodings, periodic activation functions, multi-scale or hierarchical architectures, and explicit spectral or frequency-weighted losses, all of which strengthen the network’s ability to represent and fit high-frequency content already contained in the data (tancik2020fourier; sitzmann2020implicit; liu2024mitigating; liu2025diminishing).
In the NO literature, lack of discretization-invariance (lack of DI) and DME formalize the problem that the same operator network can produce inconsistent outputs on different grids (bartolucci2023representation; gao2025discretization; lanthaler2024discretization). These works analyze aliasing and discretization error to propose alias-free spectral parameterizations, multi-grid training, and carefully designed interpolation operators to enforce cross-grid consistency. More broadly, multi-resolution generation in vision and scientific ML—whether cascaded, patch-based, or training-free—expose models to multiple resolutions (tian2023resformer; he2023scalecrafter; ho2022cascaded; bar2023multidiffusion; yu2023freedom). In this sense, the first and third steps of FRL (multi-resolution training and frequency-aware loss) follow standard practices in these research directions and are not methodologically novel.
@@ 4.2 Experiment Validation
The rapid decay of high-frequency components energy in Figure 3 reveals the deeper mechanism behind the phenomenon in Figure 2: as shown by gao2025discretization, while high-resolution single-step discretization primarily introduces errors in the low-frequency portion, a more dominant mechanism during multi-step inference is that models cannot maintain frequencies above the training Nyquist frequency. This results in the pattern shown in Table 2 where, during high-resolution inference, the proportion of low-frequency error decreases while high-frequency error becomes dominant due to more severe accumulation and amplification. These experimental results validate Theorem 2, confirming that frequency components above the training resolution’s Nyquist frequency dominate the error during high-resolution inference.
@@ 6.1.1 Datasets
Fluid Simulation: Non-reacting HIT consists of 3D homogeneous isotropic turbulence simulations with a spherical $\text{O}_{2}$ core (radius $0.25L$) embedded in $\text{CH}_{4}$ (chung2022interpretable). The dataset tracks velocity components, thermodynamic variables, and species mass fractions across 98 timesteps spanning 34 microseconds. The temporal sequence is split into 70 timesteps for training, 14 for validation, and 14 for testing. Training resolution: $32^{3}$ grid; Test resolutions: $43^{3}$, $64^{3}$, and $129^{3}$. Low-resolution fields are obtained by uniformly coarsening the original $129^{3}$ DNS snapshots by factors of $2\times$, $3\times$, and $4\times$ in each spatial direction (i.e., $129^{3}\!\to\!64^{3},43^{3},32^{3}$) using spectral low-pass filtering followed by strided sampling. We choose $32^{3}$ as the training resolution because it is the coarsest grid that still preserves the main flow structures and allows stable training for all baselines, while maximizing the Nyquist gap between the training grid and the finest test grid.
Weather Forecasting: ERA5 contains global atmospheric reanalysis data across six vertical pressure levels (200–1000 hPa) (hersbach2020era5; copernicus2017era5). Physical variables include temperature $T$ (K), horizontal wind components $(u,v)$ (m/s), and geopotential height $z$ ($\text{m}^{2}/\text{s}^{2}$). The dataset spans 360 days (January 2024–January 2025) with 6-hour temporal resolution, totaling 4320 timesteps, with the final 7 days reserved for validation. Training resolution: $180\times 90\times 6$; Test resolutions: $360\times 180\times 6$, $720\times 361\times 6$, and $1440\times 721\times 6$. Here the original ERA5 fields are stored on a $1440\times 721\times 6$ grid; we construct the low-resolution training data by regridding to $180\times 90\times 6$ via area-weighted averaging.
@@ 6.1.2 Baseline Selection and Implementation Details
We selected the latest SOTA baselines for each architecture published in NeurIPS/ICLR/ICML. Fluid Simulation: Neural SPH (GNN), DeepLag (Transformer), PARCv2 (CNN), DYffusion (Diffusion), SFNO (NO), FNODE (Neural ODE), NeuralFluid (NN) (toshev2024neural; ma2024deeplag; nguyen2024parcv2; ruhling2023dyffusion; cao2024spectral; li2024fourier; li2024neuralfluid). Weather Forecasting: WeatherGFT (Transformer), PDE-CNN (CNN), ARCI (Diffusion), Graph-EFM (GNN), ClimODE (Neural ODE) (xu2024generalizing; dona2020pde; ruhling2023dyffusion; oskarsson2024probabilistic; verma2024climode). FRL-enhanced baselines follow a unified training recipe; full hyperparameter settings are provided in Appendix C.
@@ 7 Discussion
Acknowledgements
This work is supported by the National Natural Science Foundation of China (Grant No. 62372409) and the Ministry of Science and Technology of the People’s Republic of China (Grant No. 2023ZD0120704 of Project No. 2023ZD0120700).
@@ Appendix G Complete Experimental Results
In addition to the 3D scenarios presented in Section 6.1, we conducted comprehensive experiments on 2D fluid simulation and weather forecasting tasks: (1) MegaFlow2D provides large-scale 2D external flow simulations around circular and elliptical obstacles (xu2023megaflow2d). The dataset contains 2000 distinct flow configurations with approximately 900 temporal snapshots each at 0.01s intervals. We use 1600 configurations for training, 200 for validation, and 200 for testing. Physical variables include velocity fields $(u,v)$ and pressure field $p$. All simulations maintain Reynolds number $Re=300$, fluid density $\rho=1\times 10^{3}$ $\text{kg/m}^{3}$, and kinematic viscosity $\nu=1\times 10^{-3}$ $\text{m}^{2}/\text{s}$ within a $20\text{m}\times 10\text{m}$ domain. Training resolution: $64\times 32$ grid; Test resolutions: $128\times 64$, $256\times 128$, and $512\times 256$. (2) ERA5-500hPa contains global atmospheric reanalysis data at the 500 hPa pressure level (hersbach2020era5; copernicus2017era5). Physical variables include temperature $T$ (K), horizontal wind components $(u,v)$ (m/s), and geopotential height $z$ ($\text{m}^{2}/\text{s}^{2}$). The dataset spans 360 days with 6-hour temporal resolution. Training resolution: $180\times 90$; Test resolutions: $360\times 180$, $720\times 361$, and $1440\times 721$.
@@
To enable direct inference at arbitrary resolutions without interpolation-based resampling, we adapt all baseline architectures following three unified design principles. First, we eliminate resolution-dependent components by replacing fixed-size operations with their resolution-agnostic counterparts: fully convolutional layers that can process inputs of arbitrary spatial dimensions (long2015fully), dynamic graph construction that adapts to varying node counts (zheng2024survey), and variable-length token sequences in transformers that naturally handle different sequence lengths (vaswani2017attention; zhai2023bytetransformer). Second, we adopt normalized coordinate systems where spatial positions are mapped to $0,1]^{d}$ regardless of the actual resolution, ensuring that learned spatial relationships remain valid across scales. This coordinate-based representation approach has been successfully demonstrated in implicit neural representations (sitzmann2020implicit; mildenhall2020nerf), where networks learn continuous functions of normalized coordinates rather than discrete grid positions. Third, we condition the models on resolution information either implicitly through the coordinate normalization or explicitly through resolution embedding vectors, allowing the network to adapt its processing based on the sampling density. This scale-aware conditioning strategy follows principles established in multi-scale network architectures (tan2019efficientnet), enabling models to adjust their computational patterns based on the input resolution. These modifications preserve each architecture’s inductive biases (convolutions maintain translation equivariance, graph networks preserve permutation invariance (wu2020graph), and attention mechanisms maintain their global receptive field) while enabling deployment at resolutions beyond those seen during training. In summary, we adopted various architecture-specific strategies to achieve multi-resolution inference. Although these methods differ from each other and may not be elegant, they allow us to verify that Scale Anchoring is a universally present problem. More importantly, these methods enable us to validate that the scale anchoring phenomenon persists even when models operate directly at target resolutions. We confirm that this phenomenon stems from fundamental limitations in the training data’s frequency content, rather than architectural constraints.
@@
Furthermore, we compare against SOTA ZS-SR methods specifically designed for spatiotemporal forecasting: FNO and PINO for fluid simulation (li2020fourier; li2024physics), and TNO and Climate FNO for weather prediction (diab2025temporal; jiang2023efficient). These methods represent the current best practices in zero-shot super-resolution for their respective domains.